OpenStudy (wesdg1978):

How do you find the value of df^-1/dx at x = f(a)? f(x) = 5x + 10, a = 4

OpenStudy (anonymous):

first find f(a)

OpenStudy (anonymous):

then just plug it in

OpenStudy (anonymous):

f(a)=5a+10 and u know that a=4

OpenStudy (anonymous):

so f(a)=30

OpenStudy (anonymous):

Which means x=30

OpenStudy (anonymous):

Is that \[\frac{ df ^{-1} }{ dx }\]?

OpenStudy (anonymous):

Are there any parentheses around df?

OpenStudy (anonymous):

so that means u can put the f in the denominator

OpenStudy (anonymous):

and by the way wtf is f? do u mean the inverse of f(x)?

OpenStudy (wesdg1978):

not sure, I copied and pasted it straight off of the homework sheet, so I'm going out on a limb and assuming that it means inverse of f

OpenStudy (wesdg1978):

not sure what just happened there, this thing just posted all of my previous posts over again

OpenStudy (anonymous):

can u like screenshot ur hw sheet? it just says f^-1?

OpenStudy (wesdg1978):

it says what you have posted up there\[df^-1/dx \] it's number 10 and then also the same process for number 11

OpenStudy (anonymous):

ya my computer cant read that right so find the inverse of f(x). You know how to do that? Just solve x=5y+10 for y and that will be f^-1(x)

OpenStudy (anonymous):

then once u have that substitute it in the equation for f^-1(x) and we already found out that x=30

OpenStudy (anonymous):

wait a minute wtf is d

OpenStudy (anonymous):

what does the question say

OpenStudy (wesdg1978):

find the value of df^-1/dx at x = f(a) f(x) = 5x + 10, a = 4

OpenStudy (wesdg1978):

BRB